# Laser interferometry as a robust neuromorphic platform for machine learning

### v1

@@ I Introduction
A promising area being explored for neuromorphic computing, and thus for the implementation of PNNs, is the field of photonics [12, 35]. Neuromorphic computing implementations based on photonics have several advantages such as higher energy efficiency, throughput, speed and parallelization capability [12]. However, a limitation of current approaches based on photonics is scalability and difficulty of implementing low power optical nonlinearities [12]. On the scalability front quantum photonics, particularly with continuous variable (CV) cluster states, can provide a remedy as much work has been done on scaling these systems to large sizes for measurement-based quantum computing schemes [Chen2014, Yoshikawa2016, Larsen2019, Asavanant2019]. The original proposal for a CV quantum neural network [26], as well as derivative work [5], utilized optical Kerr nonlinearities which are difficult to implement experimentally. An alternative proposal utilized repeat-until-success photon number resolving measurements on ancillary modes to introduce nonlinearity into the network [6]. However, the implementation of quantum optical nonlinearities is still a challenging task as in the classical optics case. One solution that has been explored in the classical optics approach is to utilize hybrid architectures based on optical-digital conversion [50, 54, 38] or opto-electronic conversion [48, 58, 20, 4, 11, 61, Bandyopadhyay2024]. Another approach uses the intrinsic nonlinearity within coupled semiconductor nanolaser arrays [24].
However, learning nonlinearity should not be equated with physics nonlinearity. Indeed, recent breakthroughs in optical implementations [PT2024, 55, 63, 60] stemmed from the realization that optical nonlinearities in the field variables are not needed to implement learning [PT2024] if one doesn’t encode the data in them exclusively and, rather, uses the experimental parameters for that purpose: a simple optical phase shift $\theta$ imparts field transformations $\cos\theta,\sin\theta$ that are nonlinear in $\theta$ (though linear in the fields). This has recently been proposed [55] and demonstrated [63, 60] using classical linear wave scattering.
@@
It is interesting—in particular to set the stage for future extensions to the quantum regime—to deliberately adopt a quantum optical description of the optical circuit, even though it sits firmly within the classical border in this paper. Indeed, three classes of optical circuits can be defined: (i), classical linear optics (laser interferometers with photodetectors of dark current well above the single-photon level), (ii), second-order nonlinear optics yielding only two-wave mixing (quantum squeezers, e.g. undepleted-pump optical parametric oscillators) and, (iii), true three-wave mixing as well as third- and higher order nonlinear optics. While (ii) and (iii) both enter the realm of quantum optics, only (iii) involves nonlinear quantum evolution, a.k.a. non-Gaussian Wigner function gates and/or states of the quantum fields. This is why such operations—be they implemented via third-order nonlinearities or Fock state projection—were included in previous works on quantum neural networks [26, 6, 5]. Again, the gist of linear optics implementations [PT2024, 55, 63, 60] is to eschew relying solely on field encoding, and hence the need for optical nonlinearities, in favor of optical parameter encoding.
@@ II.2 Nonlinear learning with linear optics
To approximate a fully connected linear layer present in a classical neural network we use an $M\times M$ interferometer containing beamsplitters of tunable transmissivity angle $\theta$ and fixed zero phase. To maintain consistency with classical neural networks’ fully connected layers, we employ beamsplitter operations between all M(M1)/2M(M-1)/2 pairs of modes, contrasting with conventional nearest-neighbor Mach-Zehnder interferometer (MZI) approaches [42, Clements2016]. Specifically, given the mode indices $\{1,2,...,M\}$ the interferometer applies a beamsplitter operation between each pair of modes in the set of all two-element subsets denoted as
@@ IV.2 Classification
We now examine how our model performs on classification which is the primary task used for benchmarking PNNs. For this class of supervised learning problems the model learns a function that maps inputs to discrete outputs, i.e., classes [23]. For a classification problem with $C$ classes we treated the expectation value of the $Q$ quadrature of the first $C$ modes as the output logits of the model in contrast to previous CV quantum neural network proposals which used the expectation value of the photon number number observable measured using non-Gaussian photon number detection [26, 6]. We used the cross entropy loss function, as done in Refs. [38, 55, 63, Bandyopadhyay2024, 5], in particular we used the PyTorch implementation which internally normalizes the logits into probabilities by use of the softmax function [8]. Note that the use of the softmax function is needed only during model training and not during inference.
@@ IV.2.2 Multilabel Classification
In addition to the previous sections which demonstrate this ONN model can indeed solve non-trivial nonlinear learning problems we also investigate the performance of our network on more challenging and commonly used benchmarks of multilabel classification. We used three examples of multilabel classification: the Iris data set which is a commonly used benchmark [3, 2, 64, 31, 24, 61], the handwritten digits dataset considered in Refs. [53, 55] and the vowel recognition dataset which is widely used in benchmarking of physical neural networks [50, 43, 59, 34, Bandyopadhyay2024, 37]. The first two datasets are loaded via scikit-learn.
@@ V Conclusion
We have proposed a method for constructing a linear optical neural network using laser interferometry and field displacements, which is the novelty claim of this paper. We demonstrated that this classical optical model displays remarkable resilience to photon loss which is a prominent source of imperfections in modern photonic integrated circuits [Bogaerts2020]. The linear optical character of the circuit also makes it highly suitable for in situ training as there already exist methods for computing the necessary gradients of the circuit parameters using either simple intensity or homodyne measurements. All this makes it particularly appealing for on-chip integrated optics implementations. We also investigated how our network performs on a variety of machine learning problems and its robustness to variance in the gradient estimates provided by parameter shift rules.
@@
For future work one could explore how these networks can more closely approximate the expressivity of classical neural networks. For example, classical neural networks inherently have a feed forward structure in which outputs from one layer are composed with the nonlinear transformation (activation function) in the next layer leading the network learning hierarchical and distributed representations of the inputs at each layer [8]. We provide a possible extension of our approach for implementation of this feature in the appendix but leave for future work a more rigorous examination of the expressivity of our main approach and these extensions. An interesting example of such feedforward (though a nonlinear optical one, using field encoding rather than parameter encoding) can be found in Ref. [Bandyopadhyay2024]. Lastly, extensions to implementing convolutional, recurrent and transformer layers are also important next steps.
@@ Acknowledgements.
We thank Maria Schuld and Catalina Albornoz for useful discussions. OP was supported by U.S. National Science foundation grants PHY-2514971 and ECCS-2530171. JMS acknowledges support from the National Science Foundation via DMR-2204312.
@@ Appendix A Using intensity measurements
Appendix B Feedforward activation functions using linear optics
While the nonlinearity implemented in the main text is sufficient for nonlinear learning it’s function deviates somewhat from that of nonlinear activation functions in classical neural networks which are applied to the outputs of a particular layer in the network rather than the original input data. It is interesting to see how much more closely we can mimic the function of a classical neural network as this may allow for more straightforward applications of universal function approximation theorems [14, 30], as has recently been done for nonlinear resistive networks [45], to our ONN.
Here we present two activation functions that may be implemented using the phase shift nonlinearity utilized in the main text. For simplicity of the derivation we adopt a description of the $M$-mode optical system that uses the complex field amplitudes of each mode collected into the following vector
$E=\begin{pmatrix}a_{1},\dots,a_{M}\end{pmatrix}^{T}\in\mathbb{C}^{M}.$ (26)
Under this description single mode phase shifts are defined as [22]
$R(\vec{\phi})=\begin{pmatrix}e^{i\phi_{1}}&0&\cdots&0\\ 0&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&e^{i\phi_{M}}\end{pmatrix}$ (27)
where
$\vec{\phi}=\begin{pmatrix}\phi_{1},\dots,\phi_{M}\end{pmatrix}\in\mathbb{R}^{M}$ (28)
and under which the complex field amplitudes transform as
$E\mapsto R(\vec{\phi})E.$ (29)
B.1 Complex exponential activation
Suppose $E^{\prime}$ is the resulting vector of field amplitudes after the interferometer and displacement operations from the circuit in the main text are applied in layer $j$. A straightforward implementation of a feedforward nonlinear activation function may be realized by performing intensity measurements of the light in each mode $|E^{\prime}_{i}|^{2}$ (or homodyne measurements) and feeding these measurement results as inputs to a phase shift operation $R(|E^{\prime}_{1}|^{2},\cdots,|E^{\prime}_{M}|^{2})$. This phase shift operation may then be applied to either $E^{\prime}$ or another set of optical fields where $E=[1,\dots,1]^{T}$. The latter case implements the nonlinear activation function
$\textrm{cis}(|E^{\prime}|^{2})$ (30)
where $\textrm{cis}(x)=\cos(x)+i\sin(x)=e^{ix}$ which more closely resembles the output of a classical neural networks activation function.
B.2 Gated linear units
The gated linear unit (GLU) activation function [15], which is used in classical neural networks and whose variants such as SwiGLU are commonly used in transformer based large language models [49, 36, 19, 62, 57] is also interesting to study. The GLU activation function, given an input vector $x\in\mathbb{R}^{M}$, real matrices $W,V\in\mathbb{R}^{M\times N}$ and real vectors $b,c\in\mathbb{R}^{N}$, is defined as [15, 49]
$\textrm{GLU}(x,W,V,b,c)=\sigma(xW+b)\odot(xV+c)$ (31)
where $\sigma(\cdot)$ denotes element-wise application of the sigmoid function and $\odot$ denotes the Hadamard product i.e. element-wise multiplication.
We can define an activation function analogous to the GLU function based on this phase shift operation as
$\textrm{CisGLU}(E,W,V,b,c)=R(|E^{\prime}_{1}|^{2},\dots,|E^{\prime}_{M}|^{2})(VE+c)$ (32)
where
$E^{\prime}=WE+b,$ (33)
$W,V\in\mathbb{C}^{M\times M}$, $b,c\in\mathbb{C}^{M}$, and we use Cis in reference to the complex exponential. The activation function essentially consists of feeding the field intensities of $E^{\prime}$ into phase shift operations which are then applied to the field $VE+c$. Using the fact that $R(\vec{\phi})$ is diagonal Eq. 32 can equivalently be written in a form more closely resembling Eq. 31 as
$\textrm{CisGLU}(E,W,V,b,c)$ $=\textrm{cis}(|E^{\prime}|^{2})\odot(VE+c)$ (34)
$=\begin{pmatrix}e^{i|E^{\prime}_{1}|^{2}},\dots,e^{i|E^{\prime}_{M}|^{2}}\end{pmatrix}^{T}\odot(VE+c).$ (35)
One can envisage $W,V$ may be implemented using interferometers and optical amplifiers/attenuators [50, 26], $c,d$ may be implemented using field displacements and the values $|E_{i}|^{2}$ may be extracted via intensity measurements of the optical field.
We thus have a method of implementing an ONN with a nonlinear activation function, along with the necessary linear transformations, that comprise a simple fully connected neural network using laser interferometry. We leave the application of universal approximation theorems to this ONN to future work.

### v2

@@ I Introduction
A promising area being explored for neuromorphic computing, and thus for the implementation of PNNs, is the field of photonics [12, 34]. Neuromorphic computing implementations based on photonics have several advantages such as higher energy efficiency, throughput, speed and parallelization capability [12]. However, a limitation of current approaches based on photonics is scalability and difficulty of implementing low power optical nonlinearities [12]. On the scalability front quantum photonics, particularly with continuous variable (CV) cluster states, can provide a remedy as much work has been done on scaling these systems to large sizes for measurement-based quantum computing schemes [Chen2014, Yoshikawa2016, Larsen2019, Asavanant2019]. The original proposal for a CV quantum neural network [25], as well as derivative work [5], utilized optical Kerr nonlinearities which are difficult to implement experimentally. An alternative proposal utilized repeat-until-success photon number resolving measurements on ancillary modes to introduce nonlinearity into the network [6]. However, the implementation of quantum optical nonlinearities is still a challenging task as in the classical optics case. One solution that has been explored in the classical optics approach is to utilize hybrid architectures based on optical-digital conversion [47, 51, 37] or opto-electronic conversion [46, 54, 18, 4, 11, 57, Bandyopadhyay2024]. Another approach uses the intrinsic nonlinearity within coupled semiconductor nanolaser arrays [23].
However, learning nonlinearity should not be equated with physics nonlinearity. Indeed, recent breakthroughs in optical implementations [PT2024, 52, 58, 56] stemmed from the realization that optical nonlinearities in the field variables are not needed to implement learning [PT2024] if one doesn’t encode the data in them exclusively and, rather, uses the experimental parameters for that purpose: a simple optical phase shift $\theta$ imparts field transformations $\cos\theta,\sin\theta$ that are nonlinear in $\theta$ (though linear in the fields). This has recently been proposed [52] and demonstrated [58, 56] using classical linear wave scattering.
@@
It is interesting—in particular to set the stage for future extensions to the quantum regime—to deliberately adopt a quantum optical description of the optical circuit, even though it sits firmly within the classical border in this paper. Indeed, three classes of optical circuits can be defined: (i), classical linear optics (laser interferometers with photodetectors of dark current well above the single-photon level), (ii), second-order nonlinear optics yielding only two-wave mixing (quantum squeezers, e.g. undepleted-pump optical parametric oscillators) and, (iii), true three-wave mixing as well as third- and higher order nonlinear optics. While (ii) and (iii) both enter the realm of quantum optics, only (iii) involves nonlinear quantum evolution, a.k.a. non-Gaussian Wigner function gates and/or states of the quantum fields. This is why such operations—be they implemented via third-order nonlinearities or Fock state projection—were included in previous works on quantum neural networks [25, 6, 5]. Again, the gist of linear optics implementations [PT2024, 52, 58, 56] is to eschew relying solely on field encoding, and hence the need for optical nonlinearities, in favor of optical parameter encoding.
@@ II.2 Nonlinear learning with linear optics
To approximate a fully connected linear layer present in a classical neural network we use an $M\times M$ interferometer containing beamsplitters of tunable transmissivity angle $\theta$ and fixed zero phase. To maintain consistency with classical neural networks’ fully connected layers, we employ beamsplitter operations between all M(M1)/2M(M-1)/2 pairs of modes, contrasting with conventional nearest-neighbor Mach-Zehnder interferometer (MZI) approaches [41, Clements2016]. Specifically, given the mode indices $\{1,2,...,M\}$ the interferometer applies a beamsplitter operation between each pair of modes in the set of all two-element subsets denoted as
@@ IV.2 Classification
We now examine how our model performs on classification which is the primary task used for benchmarking PNNs. For this class of supervised learning problems the model learns a function that maps inputs to discrete outputs, i.e., classes [22]. For a classification problem with $C$ classes we treated the expectation value of the $Q$ quadrature of the first $C$ modes as the output logits of the model in contrast to previous CV quantum neural network proposals which used the expectation value of the photon number number observable measured using non-Gaussian photon number detection [25, 6]. We used the cross entropy loss function, as done in Refs. [37, 52, 58, Bandyopadhyay2024, 5], in particular we used the PyTorch implementation which internally normalizes the logits into probabilities by use of the softmax function [8]. Note that the use of the softmax function is needed only during model training and not during inference.
@@ IV.2.2 Multilabel Classification
In addition to the previous sections which demonstrate this ONN model can indeed solve non-trivial nonlinear learning problems we also investigate the performance of our network on more challenging and commonly used benchmarks of multilabel classification. We used three examples of multilabel classification: the Iris data set which is a commonly used benchmark [3, 2, 59, 30, 23, 57], the handwritten digits dataset considered in Refs. [50, 52] and the vowel recognition dataset which is widely used in benchmarking of physical neural networks [47, 42, 21, 55, 33, Bandyopadhyay2024, 36]. The first two datasets are loaded via scikit-learn.
@@ V Conclusion
We have proposed a method for constructing a linear optical neural network using laser interferometry and field displacements, which is the novelty claim of this paper. We demonstrated that this classical optical model displays remarkable resilience to photon loss which is a prominent source of imperfections in modern photonic integrated circuits [Bogaerts2020]. The linear optical character of the circuit also makes it highly suitable for in situ training using various existing training protocols based on parameter shift rules [45], physical backpropagation [20], physics-aware training [55] or forward-only training [33, 35, Bandyopadhyay2024, 57]. All this makes it particularly appealing for on-chip integrated optics implementations. We also investigated how our network performs on a variety of machine learning problems and its robustness to variance in the gradient estimates provided by parameter shift rules.
@@
For future work one could explore how these networks can more closely approximate the expressivity of classical neural networks. For example, classical neural networks inherently have a feed forward structure in which outputs from one layer are composed with the nonlinear transformation (activation function) in the next layer leading to the network learning hierarchical and distributed representations of the inputs at each layer [8]. An interesting example of such feedforward (though a nonlinear optical one, using field encoding rather than parameter encoding) can be found in Ref. [Bandyopadhyay2024]. Furthermore, the existence of universal function approximation theorems for classical neural networks [14, 29] provides a strong theoretical basis for their expressivity and possible extensions of these theorems to our work merits further investigation. Lastly, extensions to implementing convolutional, recurrent and transformer layers are also important next steps.
@@ Acknowledgements.
We thank Benjamin Scellier, Maria Schuld and Catalina Albornoz for useful discussions. OP was supported by U.S. National Science foundation grants PHY-2514971 and ECCS-2530171. JMS acknowledges support from the National Science Foundation via DMR-2204312.